OpenStudy (anonymous):

Use substitution to solve the equation (Steps would help, thank you!!). 11x + 4y = -17 -6 + y = 22

OpenStudy (anonymous):

From the second equation y = 22 +6 = 28. Substitute this in equation 1 to get x = -129/11

OpenStudy (campbell_st):

well If the 2nd equation is correct y = 28 after rewriting... so sub y = 28 into the 1st equation 11x + 2(28) = -17 etc